OpenStudy (anonymous):

Choose the equation of the vertical line passing through the point (-4, 2)

OpenStudy (jdoe0001):

ok, do you know what the slope of a vertical line is?

OpenStudy (anonymous):

noo

OpenStudy (jdoe0001):

what is a slope formula anyway?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well, that'd be the slope-intercept form, so, how do we find the slope of any 2 points though?

OpenStudy (anonymous):

rise over run

OpenStudy (jdoe0001):

right, so, rise will be \(y_2 - y_1\)

OpenStudy (jdoe0001):

let's see a vertical line

OpenStudy (jdoe0001):

you have (-4, 2) so, when x = -4, y = 2 what about if x is ... wait.... is a vertical line, "x" can't be anything than just -4 but "y" can be anything 1,000, 0000 and so on

OpenStudy (jdoe0001):

so, the "rise" can be a number from 0- \(\infty but "run" is always -4, -4-4, is well "0", anyhow, the slope is undefined so, the line equation will be x = -4 just a fixed straight line, doesn't matter what "y" is, "x" will always be equal -4

OpenStudy (anonymous):

so the answer is x=-4 ?

OpenStudy (jdoe0001):

yes, all vertical lines only have 1 "x" variable component, and no "y", because it doesn't matter how big "y" gets, "x" will always be the same number

OpenStudy (anonymous):

Thanks !!!! :) help me with a few more ??

OpenStudy (jdoe0001):

sure, just post in the channel, so we all can see and help and revise each other :)

OpenStudy (anonymous):

okay :)

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = x + 2 @jdoe0001

OpenStudy (jdoe0001):

well, I meant post in the channel, as in the main channel

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

I see it :)